OpenStudy (anonymous):

If p(x) = 2x2 – 4x and q(x) = x – 3, what is (p ºq)(x)

OpenStudy (anonymous):

\[p(x)=2x^2-4x,q(x)=x-3\] and you want \[p\circ q(x)\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

SOLOMON MY BOY :))

OpenStudy (solomonzelman):

(pº q)(x) is same as p(q(x)). you take the function q(x), and plug it instead of x into the f(x).

OpenStudy (solomonzelman):

p(x) = 2x² – 4x and q(x) = x – 3 p(q(x)) = 2(q(x))²-4(q(x)) p(q(x)) = 2(x-3)²-4(x-3)

OpenStudy (solomonzelman):

then simplify

OpenStudy (anonymous):

2x2−16x+30?

OpenStudy (solomonzelman):

yes, correct!

OpenStudy (anonymous):

yes ty !!

OpenStudy (solomonzelman):

yw